# Saving SWE-Bench: A Benchmark Mutation Approach for Realistic Agent Evaluation

### v1

@@
Abstract
Current benchmarks for evaluating software engineering agents, such as SWE-Bench Verified, are predominantly derived from GitHub issues and fail to accurately reflect how developers interact with chat-based coding assistants in integrated development environments (IDEs). We posit that this mismatch leads to a systematic overestimation of agent’s capabilities in real-world scenarios, especially bug fixing. We introduce a novel benchmarking framework that transforms existing formal benchmarks into realistic user queries through systematic analysis of developer interaction patterns with chat-based agents. Our methodology is flexible and can be easily extended to existing benchmarks. In this paper, we apply our testing framework to SWE-Bench Verified, the TypeScript subset of Multi-SWE-Bench and a private benchmark, SWE-Bench C# and transform formal GitHub issue descriptions into realistic user-style queries based on telemetry analysis of a popular chat-based agent interactions. Our findings reveal that existing benchmarks significantly overestimate agent capabilities for some models by 50% over baseline performance for public benchmarks and 10-16% for our internal benchmark. This work establishes a new paradigm for evaluating interactive chat-based software engineering agents through benchmark mutation techniques.
I Introduction
The emergence of AI-powered Software engineering agents have transformed the developer landscape. Interactive chat-based agents, including Claude Code [2], VSCode Agent [11], represent a paradigm shift from traditional fully autonomous agents like GitHub Copilot Agent [5] that process complete problem specifications independently. Unlike their autonomous counterparts, interactive agents like Claude Code engage in iterative, contextual conversations with developers and solving the problem slowly over the course of the conversation. Therefore, evaluating them requires methodologies that capture the nuances of human-chat interactions in software development.
Current evaluation benchmarks for software engineering agents, particularly SWE-Bench Verified [12], suffer from some fundamental limitations when it comes to their effectiveness in assessing interactive chat-based agents. Firstly, these benchmarks are constructed from GitHub issues, which often contain detailed and well thought out problem descriptions that differ substantially from the concise, informal queries typical of IDE chat interactions. Secondly, the public availability and extensive study of these benchmarks has led to model overfitting [10], where systems perform really well on public benchmarks but fail to generalize in the same way to non-public benchmarks.
@@ I Introduction
2. Benchmark Mutation Methodology: We introduce a systematic approach for transforming formal benchmark problems into realistic initial user queries. As a way of testing our methodology on real-world benchmarks, we create SWE-Bench Verified-Mutated and SWE-Bench C#-Mutated datasets. We show how our approach is highly extensible and can be applied to almost any bug-fixing benchmark.
3. Empirical Agent Evaluation: We provide the first comprehensive evaluation of the open-source coding agent, OpenHands [18], on both original and mutated benchmarks, revealing significant performance gaps and highlighting the importance of realistic evaluation scenarios.
4. Code Artifacts: We incorporate our approach into the official SWE-Bench GitHub repo and share our working implementation consisting of all the prompts we used and mutations generated for the SWE-Bench Verified dataset111Implementation of SWE-Bench-Mutated available at: https://github.com/bstee615/SWE-Bench-Mutated. We hope that this allows the community to easily replicate our results and adopt this technique in future agent evaluations.
@@
II Background and Related Work
@@
II-A Software Engineering Benchmarks
@@
SWE-bench benchmark [8] marked a turning point in software engineering agent evaluation by moving beyond isolated function / file-level tasks to real-world repo-level problems drawn from GitHub issues. This benchmark contained tasks of higher complexity than prior benchmarks, requiring agents to navigate and understand repository context, coordinate changes across multiple files, and produce patches that pass existing as well as hidden test cases. It’s successor SWE-Bench Verified [12] addressed concerns with the quality and contamination, and provided a human-verified set of 500 examples to evaluate agents against. However, problems still remain in the benchmark. Studies like SWE-Bench+ [1] have revealed issues like solution leakage and weak test suites.
@@
SWE-bench has been expanded to new domains. SWE-Bench-Multimodal [21] introduces JavaScript tasks with visual elements, while Multi-SWE-Bench [23] targets 7 programming languages and provides expert-annotated examples. Similarly, SWE-Poly-Bench [14] also targets languages beyond Python. Zhang et al. introduce SWE-Bench-Live [24], which is a continuously updating benchmark spanning a large set of repositories that’s meant to combat model overfitting and data contamination issues. While these benchmarks have advanced the area of fixing functional bugs with coding agents substantially, our work reveals a fundamental limitation they all share i.e. their reliance on formal GitHub issues fails to capture how developers actually communicate with chat-based agents in IDE environments. Our mutation methodology uniquely addresses this gap by transforming these formal benchmark problems into more realistic user queries, enabling more accurate assessment of agent performance in real-world scenarios.
II-B Software Engineering Agents
@@
Recent work has developed sophisticated frameworks for autonomous software engineering. SWE-Agent [20] introduced us to the Agent-Computer Interface (ACI) design principles and achieved 12.5% on SWE-Bench benchmark. OpenHands [18] provides an open-source platform where agents write code, interact with the command-line, and browse the web like human developers. Interestingly, Xia et al. proposed [19] that uses a simple three-phase process of localization, repair, and validate and is able to solve 32% of the problems in SWE-Bench-Lite without a complex agent architecture unlike its predecessors.
@@
Building on reasoning capabilites of LLMs, ReAct [22] interleaves reasoning traces and task-specific actions, while Reflexion [17] employs reinforcement through verbal feedback by getting agent to reflect on task feedback signals. Studies like Toolformer [15] explore the ability of LMs to teach themselves how and when to use tools through self-supervision. In a similar vein, Gorilla [13] finetune smaller LMs to use tools and surpass the capabilities of existing state of the art LLMs.
@@
II-C Developer Behaviour & Chat-based Assistants
Several studies have been conducted on how developers use chat-based assistants. Ziegler et al. [28] measure the impact of GitHub copilot on developer productivity. Imai et al. [6] conduct an empirical study to investigate the effectiveness of pair-programming with Copilot in comparison to human-human pair programming. Scholl et al. [16] study how novice programmers use ChatGPT for programming exercises. McNutt et al. investigate the potential of code-assistants in Jupyter notebooks and examined design considerations for interactive environments. Li et al. [9] analyzed characteristics of ChatGPT usage by developers and found that developers tend to prefer short and task-focused interactions with the chat. Our work builds directly upon these studies of developer behavior with chat-based agents by conducting the first comprehensive analysis specifically focused on how developers communicate bugs to these systems. By connecting developer behavior studies directly to benchmark design, we establish a new paradigm for evaluation that better reflects the actual conditions under which coding agents operate.
II-D Data Contamination & Mutation-Based Benchmarks
Data contamination is a prevalent issue when it comes to public benchmarks. Approaches like LiveCodeBench [7] and SWE-Bench-Live [24] provide continuously updating benchmarks to mitigate the impacts of data contamination in static benchmarks. DyVal [26] dynamically generates evaluation sets using directed acyclic graphs (DAGs). LLMEval-3 [25] provides a framework for dynamically evaluating LLMs by building on a proprietary bank of graduate-level questions that are dynamically sampled for each evaluation run.
DyCodeEval [4] provides a dynamic benchmarking approach that generates diverse problem sets to ensure consistent and reliable evaluations when evaluating LLMs under contamination risks. Zhu et al. [27] propose Inference-Time-Decontamination (ITD) to address data contamination by detecting and rewriting leaked samples without altering problem difficulties. Our benchmark mutation methodology complements these contamination-mitigation approaches while addressing a different fundamental issue i.e. the gap between how developers communicate in GitHub issues vs a chat-based agent. While previous work has focused primarily on preventing model overfitting through continuous benchmark updating or dynamic generation, our approach transforms existing benchmarks to better match real-world usage patterns. This not only helps combat overfitting, as shown by the substantial performance differences between public and private benchmarks in our results, but also enables more accurate assessment of how agents perform under realistic conditions.
III Methodology
III-A Telemetry Analysis and Pattern Extraction
@@
III-A1 Data Collection and Labeling
@@
1.
Extract High-Level Categories in Sample: The first stage involved generating high-level categories for software engineering tasks. We presented the LLM with a sample of 500 user queries and use the prompt shown in Figure 8 in the Appendix to get the LLM to generate meaningful task categories. This yielded categories in Figure 2.
- 2.
@@
| API Integration & Data Processing | Bug Fixing & Error Resolution |
@@
| Code Refactoring & Code Quality | Code Search, Analysis, & Documentation |
@@
| Configuration & Dependency Management | DevOps, CI/CD, & Infrastructure |
@@
| Feature Implementation & Enhancement | Project Setup, Build, & Deployment |
@@
| Software Testing & Test Automation | User Interface (UI/UX) Design & Frontend |
@@
III-A2 Understanding Bug-Fix User Queries
To better understand the characteristics of the bug-fixing subset of user queries, we compare them against problems in various bug-fixing benchmarks for coding agents. Figure 3 shows a side by side comparison of word counts in the problem statements of public bug-fixing benchmarks such as SWE-Bench Verified [12], SWE-PolyBench [14] and Multi-SWE-bench [23] against bug-fixing user queries in telemetry. We see a dramatic difference in query length distributions. We can see that user queries tend to be lot more succinct than benchmarks, being in the range of 10-30 words, while problems from benchmarks like SWE-Bench typically contain 100 words.
To further demonstrate the differences between the communication styles of users, we measure how often benchmarks and user queries from telemetry contain various key elements that make a bug report useful, such as file-path(s), reproduction code, test case(s), expected vs actual behavior and environment / configuration details, using prompt shown in Figure 10 in the Appendix with an LLM. Figure 4 shows a comparison of how often these elements show up in the queries contained within benchmarks, containing primarily GitHub issues, and user queries. We can see that users tend to communicate more with the error stack and file-paths compared to any other element in agent queries. In stark contrast, GitHub issues contain other elements like reproduction code, expected behavior, environment details as well. These features are largely missing from user queries. Our analysis reflects chat users’ preference for sharing more targeted information like file-paths, error stacks, rather than providing exhaustive details upfront as typically seen in GitHub issues.
We also conduct a deeper analysis to identify specific communication patterns within these bug-fixing queries to understand how users communicate bugs to a chat-based agent. To systematically identify communication patterns, we employed LLMs with the template extraction prompt shown in Figure 11 in the Appendix.
@@ III-B Mutation Methodology
III-B1 Mutation Algorithm
We use LLMs for applying the mutation with the prompt shown in Figure 12 in the Appendix. The prompt takes as inputs: a problem description from an existing benchmark, the corresponding code patch that fixes the problem, and the set of communication templates extracted from telemetry analysis in Section III-A2. The algorithm asks the model to apply multiple applicable templates to generate a diverse and realistic set of variants for each problem. The reason why our mutation prompt incorporates patch information is to enable realistic references to specific files, functions, and line numbers that users often include in their queries.
@@ III-B1 Mutation Algorithm
III-B2 Mutation Example From SWE-Bench
@@
III-B3 Visualizing Mutated Queries Against Telemetry
Figure 7 shows a 3d visualization where we plot the mutations generated by our approach for SWE-Bench C# dataset, as well as the original benchmark, against our set of telemetry queries. All the queries were first embedded using text-embedding-3-large model, followed by PCA to reduce the dimensionality of the embeddings to be plotted in a 3d figure. We can see that the embeddings corresponding to SWE-Bench C# (red) and the telemetry queries (blue) make two separate clouds, with little to no overlap. However, the cloud corresponding to SWE-Bench C#-Mutated (green) has a lot more overlap with telemetry query cloud than the original benchmark. This shows that our mutation approach successfully transforms benchmark queries to better match the characteristics of actual developer interactions. While the original benchmark queries form a distinct cluster separated from real-world usage patterns, our mutation strategy shifts the distribution toward more realistic user queries. This visualization provides empirical evidence that our approach produces queries that more closely resemble how developers naturally communicate with chat-based agents in practice, validating the core premise of our benchmark mutation methodology.
IV Experimental Setup

### v3

@@
Abstract.
Current benchmarks for evaluating software engineering agents, such as SWE-Bench Verified, are predominantly derived from GitHub issues and fail to accurately reflect how developers interact with chat-based coding assistants in integrated development environments (IDEs). We posit that this mismatch leads to a systematic overestimation of agent’s capabilities in real-world scenarios, especially bug fixing. We introduce a novel benchmarking framework that transforms existing formal benchmarks into realistic user queries through systematic analysis of developer interaction patterns with chat-based agents. Our methodology is flexible and can be easily extended to existing benchmarks. In this paper, we apply our testing framework to SWE-Bench Verified, the TypeScript subset of Multi-SWE-Bench and a private benchmark, SWE-Bench C# and transform formal GitHub issue descriptions into realistic user-style queries based on telemetry analysis of a popular chat-based agent interactions. Our findings reveal that existing benchmarks significantly overestimate agent capabilities for some models by 50% over baseline performance for public benchmarks and 10-16% for our internal benchmark. This work establishes a new paradigm for evaluating interactive chat-based software engineering agents through benchmark mutation techniques. Our code is available at: https://anonymous.4open.science/r/SWE-Bench-Mutated-ED8F
1. Introduction
The emergence of AI-powered Software engineering agents have transformed the developer landscape. Interactive chat-based agents, including Claude Code (Anthropic, 2024), VSCode Agent (Microsoft, 2025), represent a paradigm shift from traditional fully autonomous agents like GitHub Copilot Agent (GitHub, 2024) that process complete problem specifications independently. Unlike their autonomous counterparts, interactive agents like Claude Code engage in iterative, contextual conversations with developers and solving the problem slowly over the course of the conversation. Therefore, evaluating them requires methodologies that capture the nuances of human-chat interactions in software development.
Current evaluation benchmarks for software engineering agents, particularly SWE-Bench Verified (OpenAI, 2024), suffer from some fundamental limitations when it comes to their effectiveness in assessing interactive chat-based agents. Firstly, these benchmarks are constructed from GitHub issues, which often contain detailed and well thought out problem descriptions that differ substantially from the concise, informal queries typical of IDE chat interactions. Secondly, the public availability and extensive study of these benchmarks has led to model overfitting (Liang et al., 2025), where systems perform really well on public benchmarks but fail to generalize in the same way to non-public benchmarks.
@@ 1. Introduction
2. Benchmark Mutation Methodology: We introduce a systematic approach for transforming formal benchmark problems into realistic initial user queries. As a way of testing our methodology on real-world benchmarks, we create SWE-Bench Verified-Mutated, Multi-SWE-Bench (TypeScript)-Mutated and SWE-Bench C#-Mutated datasets. We show how our approach is highly extensible and can be applied to almost any bug-fixing benchmark.
3. Empirical Agent Evaluation: We provide the first comprehensive evaluation of the open-source coding agent, OpenHands (Wang et al., 2024), on both original and mutated benchmarks, revealing significant performance gaps and highlighting the importance of realistic evaluation scenarios.
4. Code Artifacts: We incorporate our approach into the official SWE-Bench GitHub repo and share our working implementation consisting of all the prompts we used and mutations generated for the SWE-Bench Verified dataset111Implementation of SWE-Bench-Mutated available at: https://anonymous.4open.science/r/SWE-Bench-Mutated-ED8F. We hope that this allows the community to easily replicate our results and adopt this technique in future agent evaluations.
@@
2. Background and Related Work
@@
2.1. Software Engineering Benchmarks
@@
SWE-bench benchmark (Jimenez et al., 2024) marked a turning point in software engineering agent evaluation by moving beyond isolated function / file-level tasks to real-world repo-level problems drawn from GitHub issues. This benchmark contained tasks of higher complexity than prior benchmarks, requiring agents to navigate and understand repository context, coordinate changes across multiple files, and produce patches that pass existing as well as hidden test cases. It’s successor SWE-Bench Verified (OpenAI, 2024) addressed concerns with the quality and contamination, and provided a human-verified set of 500 examples to evaluate agents against. However, problems still remain in the benchmark. Studies like SWE-Bench+ (Aleithan et al., 2024) have revealed issues like solution leakage and weak test suites.
@@
SWE-bench has been expanded to new domains. SWE-Bench-Multimodal (Yang et al., 2024b) introduces JavaScript tasks with visual elements, while Multi-SWE-Bench (Zan et al., 2025) targets 7 programming languages and provides expert-annotated examples. Similarly, SWE-Poly-Bench (Rashid et al., 2025) also targets languages beyond Python. Zhang et al. introduce SWE-Bench-Live (Zhang et al., 2025a), which is a continuously updating benchmark spanning a large set of repositories that’s meant to combat model overfitting and data contamination issues. While these benchmarks have advanced the area of fixing functional bugs with coding agents substantially, our work reveals a fundamental limitation they all share i.e. their reliance on formal GitHub issues fails to capture how developers actually communicate with chat-based agents in IDE environments. Our mutation methodology uniquely addresses this gap by transforming these formal benchmark problems into more realistic user queries, enabling more accurate assessment of agent performance in real-world scenarios.
2.2. Software Engineering Agents
@@
Recent work has developed sophisticated frameworks for autonomous software engineering. SWE-Agent (Yang et al., 2024a) introduced us to the Agent-Computer Interface (ACI) design principles and achieved 12.5% on SWE-Bench benchmark. OpenHands (Wang et al., 2024) provides an open-source platform where agents write code, interact with the command-line, and browse the web like human developers. Interestingly, Xia et al. proposed (Xia et al., 2024) that uses a simple three-phase process of localization, repair, and validate and is able to solve 32% of the problems in SWE-Bench-Lite without a complex agent architecture unlike its predecessors.
@@
Building on reasoning capabilites of LLMs, ReAct (Yao et al., 2023) interleaves reasoning traces and task-specific actions, while Reflexion (Shinn et al., 2023) employs reinforcement through verbal feedback by getting agent to reflect on task feedback signals. Studies like Toolformer (Schick et al., 2023) explore the ability of LMs to teach themselves how and when to use tools through self-supervision. In a similar vein, Gorilla (Patil et al., 2023) finetune smaller LMs to use tools and surpass the capabilities of existing state of the art LLMs.
@@
2.3. Developer Behaviour & Chat-based Assistants
Several studies have been conducted on how developers use chat-based assistants. Ziegler et al. (Ziegler et al., 2024) measure the impact of GitHub copilot on developer productivity. Imai et al. (Imai, 2022) conduct an empirical study to investigate the effectiveness of pair-programming with Copilot in comparison to human-human pair programming. Scholl et al. (Scholl and Kiesler, 2024) study how novice programmers use ChatGPT for programming exercises. McNutt et al. investigate the potential of code-assistants in Jupyter notebooks and examined design considerations for interactive environments. Li et al. (Li et al., 2025) analyzed characteristics of ChatGPT usage by developers and found that developers tend to prefer short and task-focused interactions with the chat. Our work builds directly upon these studies of developer behavior with chat-based agents by conducting the first comprehensive analysis specifically focused on how developers communicate bugs to these systems. By connecting developer behavior studies directly to benchmark design, we establish a new paradigm for evaluation that better reflects the actual conditions under which coding agents operate.
2.4. Data Contamination & Mutation-Based Benchmarks
Data contamination is a prevalent issue when it comes to public benchmarks. Approaches like LiveCodeBench (Jain et al., 2024) and SWE-Bench-Live (Zhang et al., 2025a) provide continuously updating benchmarks to mitigate the impacts of data contamination in static benchmarks. DyVal (Zhu et al., 2024a) dynamically generates evaluation sets using directed acyclic graphs (DAGs). LLMEval-3 (Zhang et al., 2025b) provides a framework for dynamically evaluating LLMs by building on a proprietary bank of graduate-level questions that are dynamically sampled for each evaluation run.
DyCodeEval (Chen et al., 2025) provides a dynamic benchmarking approach that generates diverse problem sets to ensure consistent and reliable evaluations when evaluating LLMs under contamination risks. Zhu et al. (Zhu et al., 2024b) propose Inference-Time-Decontamination (ITD) to address data contamination by detecting and rewriting leaked samples without altering problem difficulties. Our benchmark mutation methodology complements these contamination-mitigation approaches while addressing a different fundamental issue i.e. the gap between how developers communicate in GitHub issues vs a chat-based agent. While previous work has focused primarily on preventing model overfitting through continuous benchmark updating or dynamic generation, our approach transforms existing benchmarks to better match real-world usage patterns. This not only helps combat overfitting, as shown by the substantial performance differences between public and private benchmarks in our results, but also enables more accurate assessment of how agents perform under realistic conditions.
3. Methodology
3.1. Telemetry Analysis and Pattern Extraction
@@
3.1.1. Data Collection and Labeling
@@
Extract High-Level Categories in Sample: The first stage involved generating high-level categories for software engineering tasks. We presented the LLM with a sample of 500 user queries and use the prompt shown in Figure 1 to get the LLM to generate meaningful task categories. This yielded categories in Figure 4.
- (2)
@@
| A. API Integration & Data Processing | B. Bug Fixing & Error Resolution |
@@
| C. Code Refactoring & Code Quality | D. Code Search, Analysis, & Documentation |
@@
| E. Configuration & Dependency Management | F. DevOps, CI/CD, & Infrastructure |
@@
| G. Feature Implementation & Enhancement | H. Project Setup, Build, & Deployment |
@@
| I. Software Testing & Test Automation | J. User Interface (UI/UX) Design & Frontend |
@@
3.1.2. Understanding Bug-Fix User Queries
To better understand the characteristics of the bug-fixing subset of user queries, we compare them against problems in various bug-fixing benchmarks for coding agents. Figure 5 shows a side by side comparison of word counts in the problem statements of public bug-fixing benchmarks such as SWE-Bench Verified (OpenAI, 2024), SWE-PolyBench (Rashid et al., 2025) and Multi-SWE-bench (Zan et al., 2025) against bug-fixing user queries in telemetry. We see a dramatic difference in query length distributions. We can see that user queries tend to be lot more succinct than benchmarks, being in the range of 10-30 words, while problems from benchmarks like SWE-Bench typically contain 100 words.
To further demonstrate the differences between the communication styles of users, we measure how often benchmarks and user queries from telemetry contain various key elements that make a bug report useful, such as file-path(s), reproduction code, test case(s), expected vs actual behavior and environment / configuration details, using prompt shown in Figure 7 with an LLM. Figure 6 shows a comparison of how often these elements show up in the queries contained within benchmarks, containing primarily GitHub issues, and user queries. We can see that users tend to communicate more with the error stack and file-paths compared to any other element in agent queries. In stark contrast, GitHub issues contain other elements like reproduction code, expected behavior, environment details as well. These features are largely missing from user queries. Our analysis reflects chat users’ preference for sharing more targeted information like file-paths, error stacks, rather than providing exhaustive details upfront as typically seen in GitHub issues.
We also conduct a deeper analysis to identify specific communication patterns within these bug-fixing queries to understand how users communicate bugs to a chat-based agent. To systematically identify communication patterns, we employed LLMs with the template extraction prompt shown in Figure 8.
@@ 3.2. Mutation Methodology
3.2.1. Mutation Algorithm
We use LLMs for applying the mutation with the prompt shown in Figure 11. The prompt takes as inputs: a problem description from an existing benchmark, the corresponding code patch that fixes the problem, and the set of communication templates extracted from telemetry analysis in Section 3.1.2. The algorithm asks the model to apply multiple applicable templates to generate a diverse and realistic set of variants for each problem. The reason why our mutation prompt incorporates patch information is to enable realistic references to specific files, functions, and line numbers that users often include in their queries.
@@ 3.2.1. Mutation Algorithm
3.2.2. Mutation Example From SWE-Bench
@@
3.2.3. Visualizing Mutated Queries Against Telemetry
Figure 12 shows a 3d visualization where we plot the mutations generated by our approach for SWE-Bench C# dataset, as well as the original benchmark, against our set of telemetry queries. All the queries were embedded using text-embedding-3-large model, followed by PCA to reduce the dimensionality of the embeddings to be plotted in a 3d figure. We can see that the embeddings corresponding to SWE-Bench C# (red) and the telemetry queries (blue) make two separate clouds, with little to no overlap. However, the cloud corresponding to SWE-Bench C#-Mutated (green) has a lot more overlap with telemetry query cloud than the original benchmark. This shows that our mutation approach successfully transforms benchmark queries to better match the characteristics of actual developer interactions. While the original benchmark queries form a distinct cluster separated from real-world usage patterns, our mutation strategy shifts the distribution toward more realistic user queries. This visualization provides further evidence that our approach produces queries that more closely resemble how developers naturally communicate with chat-based agents in practice, validating the core premise of our benchmark mutation methodology.
4. Experimental Setup